a production line makes cameras in a day (8 hours) 200 caneras are made sold for 327 each 20 workers attend the line and are paid 20 a hour the material cost is $58 per csmera the cost of cspital fir the plant is $5000 a day what is the multifactor productivity
Answer #1

Answer:

Multifactor productivity = output / (labour +material +capital)

Output = 200cameras * $327 = $ 65400

Labour = 20workers * $20 * 8hrs = $3200

Material = $58 * 200cameras = $11600

Capital = $ 5000

Multifactor productivity= 65400/(3200+11600+5000) . = 3.03